find general solution using variation of parameters

y" - 2y' + y = e^x/(1 + x^2)
Answer #1

Solution Given difforential equation yhl_24+y= (1+x2) Auxilary equation mlemt10 stool lllou xax 3 txab Y.(x) = 7thD Y2 Consi- 4h (х+1) У (x)= • ex. У (1-+ х) dx pix 1 , I dx х!) -an-x Hence Yp (x ) = - ex In (x²41) ot yet tanty Y = Yet 9, + Ур | Ув,